les Data Table 3: Analysis of Milk of Magnesia ge ne err Trial 1 Trial 2 Mass of milk of magnesia 1.14 g/ml Density of milk of magnesia Volume of acetic acid, initial Volume of acetic acid, final Volume of acetic acid, total Concentration of acetic acid Moles of acetic acid 0.88 M Moles of Mg(OH)2 Moles Mg(OH)2/ g milk of magnesia
Answer #1

Moles of acetic acid = Mass of acetic acid / Molecular Mass = 5 *0.88 g / 60g/mol = 0.073moles

Moles of Mg(OH)2 = 2.5 / 58.3 = 0.042

Moles of Mg(OH)2 / g milk of Magnesia = 0.042/ 2.5 = 0.017